OpenStudy (anonymous):

m-3 over 2 - m-4 over 3 = m-2 over 4

hero (hero):

\[\frac{m-3}{2} - \frac{m-4}{3} = \frac{m-2}{4}\]

hero (hero):

Multiply both sides by 12: 6(m-3) - 4(m-4) = 3(m-2)

hero (hero):

6m - 18 - 4m + 16 = 3m - 6 2m - 2 = 3m - 6 m = 8

OpenStudy (anonymous):

according to the answer key, m = 4

OpenStudy (anonymous):

2m - 3m = 1m or m 6 - 2 = 4 m = 4

hero (hero):

Yes, I added 6 to both sides, so m = 4

hero (hero):

6m - 18 - 4m + 16 = 3m - 6 2m - 2 = 3m - 6 m = 4

hero (hero):

All my steps were right the first time except for the very last step when I put m = 8 instead of m = 4

OpenStudy (anonymous):

okay. I understand this now. Thank you.